# Modular Compilation for Quantum Chiplet Architectures

### v3

@@ 2. Background and Motivation
Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet edges. Gates that are unimplementable on hardware are shown in red.
To execute a program on a target device, the logical program representation must be transformed such that it conforms to the idiosyncratic physical constraints for that particular device. This process of logical-to-physical translation represents a critical part of compilation and is present even in exotic computing technologies, e.g., quantum. In quantum computing, programs are (usually) represented as quantum circuits, and devices are subject to domain-specific physical constraints, such as decoherence time (mike_and_ike), qubit topology (the connectivity between physical qubits) (qubit_allocation; qubit_routing), and the set of basis gates (the family of operations physically implemented on the device). Thus, quantum circuit compilation is necessary to actually run quantum programs on real-world devices.
Compilation is often divided into distinct stages such as qubit allocation (layout and routing), basis translation, and optimization. The relations between each stage and the physical constraints of both monolithic and chiplet architectures are described below.
@@
Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device. It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized, technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum). The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation). Thus, algorithms often adopt various strategies to make the problem more tractable. One common method is to further divide qubit allocation into two substages: layout and routing. Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap). In other words, the qubit layout stage decides the initial position of logical qubits on the device. In contrast, qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology. Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap). In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation. Namely, an initial layout is used to perform qubit routing, which is used to generate a new initial layout.
Because qubit routing remains NP-hard (qubit_routing), another method of simplifying the problem is to assume qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration. This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph. For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality. However, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets). Thus, we could expect the solutions generated from the simplified model to be worse than a more physically accurate model, for example, in circumstances where the shortest path in terms of the quantity or depth of inserted SWAPs diverges from the shortest path in terms of fidelity, gate duration, or some combination thereof.
Basis Translation
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike). Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets. In the general case, the cross-chiplet links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low). Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors), and do not expect these heterogenous, non-universal basis gate sets, can generate incompatible outputs when targeting modular architectures.
Optimization
Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and/or parallelize quantum gates (optimization1). Through commutation and cancellation identities, it is possible to search for optimization opportunities by iteratively transforming the quantum circuit without affecting the underlying operation (optimization1). However, in general, optimization is highly time and resource intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor). As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit). One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks. Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize the negative impact on the output solution. From this perspective, modular architectures are naturally synergistic with partitioning, since cross-chiplet connections present a physically motivated boundary between operations.
3. Compilation for Modular Quantum Systems
Modular quantum architectures introduce new challenges to quantum compilation that impact the correctness of solutions and the speed of compilation. As discussed previously in §2, our model of modular quantum architectures considers the possibility that inter-chiplet links can be non-universal. This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-chiplet links). However, many compiler designs (including designs for modular architectures, see §7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets. For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors). While it is possible to create a Qiskit Backend that supports heterogeneous basis gate sets, the compiler itself does not take them into account. Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors. Moreover, compilers are called to target increasingly larger quantum devices, enabled by modularity. To address the mounting pressure of ever-growing compilation workloads, these compilers must be cognizant of classical hardware limitations and tradeoffs between solution quality and compilation latency.
To ameliorate these issues, we present a family of compilers optimized for modular quantum architectures that support chiplet links with non-universal basis gate sets, and leverage modular compilation strategies. In Section 3.1, we propose a peephole compilation pass that can be inserted into any compiler to add support for non-universal basis gate sets. This would allow any compiler to be compatible with chiplet-based devices without significant changes to their underlying implementations. Next, in Section 3.2, we introduce the stratify-elaborate quantum compiler (SEQC), which advocates for a hierarchical two-stage compilation that produces both better circuits for execution and faster compilation times.
3.1. Chiplet-awareness via Peephole Correction
The design of the peephole correction pass was inspired by the “peephole” optimizations commonly used in classical compilation (dragon_book). Peephole passes operate within a nominal peephole window and substitute its contents with a logically equivalent replacement. Notable examples in the context of quantum compilation include basis translation and elimination of self-adjoint gates. Peephole correction (Figure 1) operates on a circuit generated by qubit allocation, where it detects when non-SWAP gates are placed on non-universal inter-chiplet links (Figure 1(b)), and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links (Figure 1(c)). Additional “return” SWAPs are then inserted to return the qubits to their initial locations (Figure 1(d)). Overall, for every inter-chiplet gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates — 2 inter-chiplet and 2 intra-chiplet. Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is unaware of chiplet-specific device constraints.
For example, skipping the return SWAP operations would initially appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate. However, this hypothetical design would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit. From the perspective of modular compilation, this results in serial dependencies that prevent parallel execution of the pass. From a routing perspective, the resultant qubit permutations disrupt the post-routing layout of qubits, leading to destructive routing errors that can only be resolved by an entirely new routing pass. Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method.
3.2. The Stratify-Elaborate Quantum Compiler
We incorporate the hierarchical elements of chiplet architectures into SEQC by adopting a two-stage procedure (Figure 2). The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
3.2.1. Stratification Stage
(a) Example chiplet topology
(b) Qubit-to-subcircuit mapping
(c) Subcircuit-to-chiplet routing
(d) Inter-chiplet routing
Figure 3. Step-by-step example of inter-chiplet compilation in the Stratification stage.
In this stage, we map a quantum circuit to the inter-chiplet device topology (Figure 3(a) in our running example) by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps (Figures 3(b)3(d)). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
Qubit-to-Subcircuit Mapping
Algorithm 1 Qubit-to-Subcircuit Mapping with Sim. Annealing

### v4

@@ 2. Background and Motivation
Figure 1. Peephole correction for inter-module gates that are unimplementable on hardware (i.e., non-SWAP, shown in red).
Basis Translation
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Basis translation can be performed efficiently given the target gate set is universal (mike_and_ike), and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent modules. In the general case, the cross-module links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low). Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors), and do not expect these heterogenous, non-universal basis gate sets, can generate incompatible outputs when targeting modular architectures.
Optimization
Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and parallelize quantum gates (optimization1). Optimization is highly time and resource intensive, with the problem scaling exponentially to the number of qubits and circuit depth (qfactor). As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit). One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks. Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize the negative impact on the output solution. Modular architectures are naturally synergistic with partitioning, since cross-module connections present a physically motivated boundary between operations.
3. Compilation for Hybrid Modular Quantum Systems
While SEQC can be applied on a modular quantum system comprising hybrid QPUs, in this paper we consider the simpler case of homogeneous modules as a starting point. The reasons behind this decision are simple. First, the primary goal of this paper is to introduce and evaluate the impact of the circuit stratification compilation stage with STA and Simulated Annealing we developed, along with qubit pinning, and symbiotic swap insertion. The presence of diverse-modality QPUs with varying capabilities would dilute these results, and make it hard to separate their effect from the effects of heterogeneity. At the same time, it would make comparisons to the existing state-of-the-art hard, as, to the best of our knowledge, there are no compilers that can target multiple modalities simultaneously for different parts of the same input circuit. Thus, in order to isolate the impact of our techniques, and facilitate comparisons with existing tools, in this paper we describe and evaluate SEQC assuming homogeneous modules of the same technology. When extending SEQC to a hybrid system, one only needs to execute different elaboration stages, one for each QPU modality, but the stratification stage and the other techniques we develop will remain largely unaffected. Thus, for the remainder of this paper, we will be focusing on homogeneous QPUs based on transmons, and we leave qubit-specific stratification optimizations, and the implementation of elaboration stages for other qubit species for future work.
Modular quantum architectures introduce new challenges to quantum compilation that impact the correctness of solutions and the speed of compilation. As discussed previously in §2, our model of modular quantum architectures considers the possibility that inter-module links can be non-universal. This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-module links). However, many compiler designs (including designs for modular architectures, see §7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets. For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors). While it is possible to create a Qiskit Backend that supports heterogeneous basis gate sets, the compiler itself does not take them into account. Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors. Moreover, compilers are called to target increasingly larger quantum devices, enabled by modularity. To address the mounting pressure of ever-growing compilation workloads, these compilers must be cognizant of classical hardware limitations and tradeoffs between solution quality and compilation latency.
To ameliorate these issues, we present a family of compilers optimized for modular quantum architectures that support module links with non-universal basis gate sets, and leverage modular compilation strategies. In Section 3.1, we propose a peephole compilation pass that can be inserted into any compiler to add support for non-universal basis gate sets. This would allow any compiler to be compatible with module-based devices without significant changes to their underlying implementations. Next, in Section 3.2, we introduce the stratify-elaborate quantum compiler (SEQC), which advocates for a hierarchical two-stage compilation that produces both better circuits for execution and faster compilation times.
3.1. Module-awareness via Peephole Correction
The design of the peephole correction pass was inspired by the “peephole” optimizations commonly used in classical compilation (dragon_book). Peephole passes operate within a nominal peephole window and substitute its contents with a logically equivalent replacement. Peephole correction (Figure 1) operates on a circuit generated by qubit allocation, where it detects when non-SWAP gates are placed on non-universal inter-module links (Figure 1(b)), and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-module links (Figure 1(c)). Additional “return” SWAPs are then inserted to return the qubits to their initial locations (Figure 1(d)). Overall, for every inter-module gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates — 2 inter-module and 2 intra-module. Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is unaware of module-specific device constraints.
For example, skipping the return SWAP operations would appear to save $2$ SWAPs, and halve the cost per inter-module gate. However, it would also introduce a net qubit permutation that propagates downstream in the circuit. From the perspective of modular compilation, this creates serial dependencies that prevent parallel execution of the pass. From a routing perspective, the resultant qubit permutations disrupt the post-routing layout of qubits, causing destructive routing errors that can only be resolved by an entirely new routing pass. Overall, such a hypothetical pass would result in substantially worse output circuits compared to the peephole method.
3.2. The Stratify-Elaborate Quantum Compiler
We incorporate the hierarchical elements of module architectures into SEQC by adopting a two-stage procedure (Figure 2). The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the module level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
3.2.1. Stratification Stage
(a) Example module topology
(b) Qubit-to-subcircuit mapping
(c) Subcircuit-to-module routing
(d) Inter-module routing
Figure 3. Step-by-step example of inter-module compilation in the Stratification stage.
In this stage, we map a quantum circuit to the inter-module device topology (Figure 3(a) in our running example) by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps (Figures 3(b)3(d)). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of module allocation, we define a mapping of logical subcircuits to physical modules on the target device and route inter-module gates to conform with constraints of module connectivity.
Qubit-to-Subcircuit Mapping
Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits (numbers in Figure 3(b)) are grouped into subcircuits (lowercase letters). An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-module SWAPs needed to transfer data between modules during execution. To perform this task, we investigate two methods: Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) (ovide2024sta).
Simulated annealing is an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989). It is an iterative method to optimize a state against a given cost function. In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost. However, the algorithm may randomly also accept worse solutions to escape from local minima wells. The runtime of simulated annealing is statically determined using constant tuning parameters, thus the overall complexity is determined by the complexity of the cost and kick functions, which are typically polynomial. Additionally, since each run of a simulated annealing is randomized and independent, it is possible to run multiple trials in parallel with different initial random seeds, then pick the best result. This improves the chances of finding a (near-)optimal solution. Our simulated annealing method (Algorithm 1) starts with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined as the number of inter-module gates given a certain layout. By minimizing the number of inter-module gates, we implicitly reduce the inter-module connectivity. The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits. In doing so, we hope to move strongly interacting qubits close together, minimizing the number of cross-module gates. As input, our algorithm takes the input circuit $Circ$, the module count $N_{C}$, the module size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling factor $\beta$. The last three are constant tuning parameters across runs.
Algorithm 1 Qubit-to-Subcircuit Mapping with Simulated Annealing
1: procedure Cost($Circ,L$)
2:   $C\leftarrow 0$ $\triangleright$ Count
3:   for $g\in\textsc{TwoQubitGates}(Circ)$ do
4:     $q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
5:     $s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
6:     Increment $C$ if $s_{0}\neq s_{1}$
7:   return $C$
8: procedure Kick($D,N_{C},N_{Q},L$)
9:   $L^{\prime}\leftarrow\textsc{Copy}(L)$
10:   for $i=1,\dots,D$ do
11:     $s_{0}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\})$
12:     $s_{1}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\}-\{s_{0}\})$
13:     $q_{0},q_{1}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{0}]),\textsc{RandChoice}(L^{\prime}[s_{1}])$
14:     $L^{\prime}[s_{0}],L^{\prime}[s_{1}]\leftarrow(L^{\prime}[s_{0}]\cup\{q_{1}\})-\{q_{0}\},(L^{\prime}[s_{1}]\cup\{q_{0}\})-\{q_{1}\}$ $\triangleright$ Swap $q_{0}$ and $q_{1}$
15:   return $L^{\prime}$
16: procedure Q2SMappingAN($Circ,N_{C},N_{Q},T_{0},\alpha,\beta$)
17:   $L\leftarrow$ Initial Layout $\triangleright$ Current Solution
18:   $R\leftarrow L$ $\triangleright$ Best Solution
19:   $T\leftarrow T_{0}$ $\triangleright$ Annealing “Temperature”
20:   while $T>1$ do
21:     $D\leftarrow\textsc{Ceil}(T\times\alpha)$ $\triangleright$ Number of changes to make
22:     $L^{\prime}\leftarrow\textsc{Kick}(D,N_{C},N_{Q},L)$
23:     $\Delta\leftarrow\textsc{Cost}(Circ,L^{\prime})-\textsc{Cost}(Circ,L)$
24:     $k\leftarrow\textsc{Rand}(0,1)$ $\triangleright$ Random number in $0,1)\subset\mathbb{R}$
25:     $L\leftarrow L^{\prime}$ if $(\Delta<0)$ or $(k>\exp(\Delta/T))$
26:     $R\leftarrow L^{\prime}$ if $(\textsc{Cost}(Circ,L^{\prime})<\textsc{Cost}(Circ,R))$
27:     $T\leftarrow T\times\beta$ $\triangleright$ Decrease temperature
28:   return $R$
The STA algorithm is a qubit allocation algorithm designed for trapped-ion quantum systems. These devices often feature modular designs, where one device is composed of “traps” arranged in a (typically) line or ring topology, where each trap contains multiple qubits (ions). Qubit allocation algorithms for modular trapped-ion systems are incentivized to minimize the number of operations between modules due to associated qubit shuttling and relocation overheads (ovide2024sta). When compiling for modular transmon architectures, we are similarly incentivized to minimize the number of inter-module interactions. Thus, we adapted a component of the STA algorithm (Algorithm 2 in Ovide et al. (ovide2024sta)) that assigns qubits to ion traps to perform qubit-to-subcircuit mapping and generate an initial subcircuit-to-module mapping (Algorithm 2).
The original STA algorithm prioritizes mapping qubits that frequently interact with other qubits (quantified by $R$ in Algorithm 2), or have a large degree of direct or indirect interactions (quantified heuristically as the depth-weighted sum $T$ of interactions between every pair of qubits). These are used to derive a sequence of qubit pairs, from which STA constructs a qubit mapping. Namely, for each iteration of the STA algorithm, a pair of qubits $(q_{a},q_{b})$ is used to insert at least one qubit into the mapping, based on the shortest topological distance between physical qubits. If one of the qubits in the pair is already mapped, say $q_{a}$, then STA also recursively maps any qubits that are more strongly connected to $q_{a}$ than $q_{b}$ is. This is done in order to keep strongly connected qubits physically close together.
Algorithm 2 Qubit-to-Subcircuit Mapping with STA
1: procedure ComputeRatios($C$)
2:   $R\leftarrow$ empty map
3:   for $q\in\textsc{Qubits}(C)$ do
4:     $R[q]\leftarrow\frac{\textsc{InteractionCnt}(q,C)}{\textsc{NumQubits}(C)}$ $\triangleright$ fraction of qubits $q$ interacts with
5:   return $R$
6: procedure ComputeST($C,S$)
7:   $T\leftarrow$ empty map
8:   for $(q_{1},q_{2})\in\textsc{AllPairs}(\textsc{Qubits}(C))$ do
9:     for $(i,s)\in\textsc{enumerate}(S)$ do
10:      Add $2^{-i}$ to $T[(q_{1},q_{2})]$ if $q_{1}$ and $q_{2}$ interact in $s$
11:   return $T$
12: procedure MapQubit($q_{1},T,R,L$)
13:   $P\leftarrow$ first pair in $T$ containing $q_{1}$; $q_{2}\leftarrow$ other qubit in $P$
14:   if $\textsc{Index}(q_{2})<\textsc{Index}(q_{1})$ then
15:     $\textsc{MapQubit}(q_{2},T)$ $\triangleright$ Map $q_{2}$ if it comes first
16:   if $q_{1}\in R$ or $q_{2}\in R$ then
17:     $\textsc{PlaceQubits}(q_{1},q_{2},L)$
18:     $R.\textsc{remove}(q_{1},q_{2})$; $T.\textsc{remove}(P)$ $\triangleright$ Remove placed qubits
19: procedure Q2SMappingSTA($C$) $\triangleright$ $C$: Quantum Circuit
20:   $S\leftarrow\textsc{Layers}(C)$; $L\leftarrow$ initial layout
21:   $R\leftarrow\textsc{ComputeRatios}(C)$; $T\leftarrow\textsc{ComputeST}(S)$
22:   while $R$ is not empty do
23:     $\textsc{MapQubit}(R[0],T,R,L)$
24:   return $L$
Module Allocation
Algorithm 3 Inter-Module Routing
1: procedure FindSwaps($Q,E$)
2:   return $\textsc{FindSymbioticSwaps}(Q,E)$ if non-empty
3:   return $\textsc{FindCommensalisticSwaps}(Q,E)$ if non-empty
4:   return $\textsc{FindParasiticSwaps}(Q,E)$
5: procedure ValidGates($Q,L$)
6:   Initialize $V_{virt}$ and $V_{phys}$ to empty sets
7:   for $g\in Q$ do
8:     $H\leftarrow\{L[q]|q\in\textsc{Qubits}(g)\}$ $\triangleright$ Get module info
9:     if $|H|=1$ then $\triangleright$ All qubits on same module
10:      Add $g$ to $V_{Virt}$ and $(g,H)$ to $V_{phys}$
11:   return $V_{virt},V_{phys}$
12: procedure Step($Q,R,L,D$) $\triangleright$ Advances the “wavefront”
13:   $V_{virt},V_{phys}\leftarrow\textsc{ValidGates}(Q,L)$
14:   $R\leftarrow R+V_{phys}$
15:   for $g\in V_{virt}$ do
16:     Remove $g$ from $Q$
17:     for $c\in\textsc{Children}(g,D)$ do
18:      Add $c$ to $Q$ if $p\in R$ for all $p\in\textsc{Parents}(c,D)$
19: procedure Heuristic($S,R,Circ$)
20:   $L^{\prime}\leftarrow\textsc{Swap}(\textsc{Copy}(L),S)$; $C\leftarrow 0$
21:   for $g\in\textsc{TwoQubitGates}(Circ)$ do
22:     if $g\notin R$ then
23:      $q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
24:      $s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
25:      Increment $C$ if $s_{0}\neq s_{1}$
26:   return $C$
27: procedure InterModuleRouting($Circ,L,T$) $\triangleright$ $L$ is from Algorithm 2, $T$ is the module topology.
28:   $D\leftarrow\textsc{CircuitToDAG}(Circ)$
29:   $E\leftarrow$ All-pairs distances from module topology $T$
30:   $Q\leftarrow\textsc{FirstLayer}(DAG)$; $R\leftarrow$ empty list
31:   $\textsc{Step}(Q,R,L,D)$
32:   while $Q$ is not empty do
33:     $\mathcal{S}\leftarrow\textnormal{{FindSwaps}}(Q,E)$ $\triangleright$ SWAP Candidates
34:     $S\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{S}}\textsc{Heuristic}(s,R,Circ)$
35:     $\textsc{Swap}(L,S)$
36:     $\textsc{Step}(Q,R,L,D)$
37:   $D^{\prime}\leftarrow$ DAG built in topological order from $R$ $\triangleright$ $R$ is ordered
[…]
Module allocation describes the process of transforming a stratified circuit to be compatible with the inter-module topology of a modular device. In Figure 2, we decompose module allocation into subcircuit-to-module mapping and inter-module routing, analogous to how qubit allocation can be decomposed into qubit layout and routing. Accordingly, subcircuit-to-module mapping (Figure 3(c)), defines an initial placement of logical subcircuits (lowercase letters) to physical modules (uppercase letters). Subsequently, inter-module routing (Figure 3(d)) inserts inter-module SWAPs to account for limited cross-module gate connectivity.
In our design, we use our qubit-to-subcircuit mapping to provide an initial subcircuit-to-module mapping. Then, we extend and modify the SABRE algorithm (sabre_swap) to apply to modular architectures, placing iterations of inter-module routing alongside randomized qubit permutations to inform a final subcircuit-to-module mapping. We did explore a design that extended other components of STA to perform subcircuit-to-module mapping on transmon module systems, but it underperformed this iterative approach, and had increased compilation time, hence it was abandoned.
Inter-module routing seeks to consolidate the qubits of (non-SWAP) inter-module gates to a single module by inserting inter-module SWAPs into the stratified circuit. For a single two-qubit, inter-module gate, we represent inter-module routing as a shortest-path problem, where the two qubits of the gate correspond to module endpoints in the unweighted inter-module connectivity graph. From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or module distance, indicates the length of the SWAP chain. However, in the context of a stratified circuit, composed of multiple inter-module gates, it becomes possible to interleave SWAP chains (i.e., efficiently reorder their execution) to reduce the depth and/or total number of inter-module SWAPs. Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-module routing at compile-time. Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic. For our inter-module routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 3.
An inter-module SWAP is considered symbiotic if it directly affects two inter-module gates in a “beneficial” manner. In other words, each qubit of a symbiotic SWAP corresponds to a different inter-module gate, where the insertion of the SWAP reduces the module distance score of both gates. In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-module routing cost heuristic. Next, commensalistic SWAPs are those that do not “harm,” or increase that module distance, of any directly impacted gates. These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-module gate without affecting other inter-module gates. In certain device topologies, a SWAP acting on the qubits of two inter-module gates can also exhibit commensalistic behavior. Finally, parasitic SWAPs act on the qubits of two inter-module gates, where one gate benefits while the other is harmed. Since this variety of SWAPs is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock. For example, clearing one gate may free physical qubits to help route others.
3.2.2. Elaboration Stage
(a) Intra-module topology
example
(b) Stratified 3-qubit subcircuit
(c) Subcircuit after intra-module layout
$q_{0},q_{1},q_{2}]\Rightarrow[Q_{2},Q_{0},Q_{1}]$
(d) Inter-module SWAP lowering
(e) Intra-module routing
(f) Basis translation and optimization
Figure 4. Step-by-step example of intra-module compilation in the Elaboration stage. Gates that cannot be implemented on hardware are shown in red.
Following the stratification stage, each module can be compiled almost completely independently from one another. More precisely, the compilation process within each module is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-module SWAPs, which must be lowered from the module level to the qubit level. Once this assignment is made, these SWAPs must remain fixed, which we call “qubit pinning” and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization. To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 4 illustrates this with an example of elaboration for a single module (Figure 4(a)) and its corresponding subcircuit (Figure 4(b)).
@@ 3.2.2. Elaboration Stage
We employ a modified version of SABRE to perform qubit allocation in the elaboration stage. Specifically, we parallelize the SABRE routing algorithm and add constraints to maintain correctness when applied to a stratified quantum circuit with inter-module gates. We also explored a layout pass that used portions of the STA algorithm, but the marginal benefits to solution quality were not worth the $5-6\times$ slower elaboration time.
The placement of inter-module gates is a non-parallelizable task (due to cross-module interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across modules. To account for this, we introduce a serial stage (Figure 4(d)) between the parallel layout and routing passes that greedily places gates on the nearest valid inter-module edge. We choose a greedy policy because inter-module gates have more restrictions on valid placements compared to intra-module gates, while also experiencing substantially higher error rates (smith_2022_chiplets). Once the positions of the inter-module gates are settled, they are pinned, imposing an additional constraint on qubit routing. To softly encourage this behavior in SABRE, we extend its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-module gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-module SWAPs. Thus, we need to impose hard restrictions to ensure correctness. To address this we establish two disjoint gate sets: one gate set that operates strictly within a module (the traditional gate set for a monolithic processor), and another that is the only gate set permitted to operate across modules (and, specifically for our example implementation, only implements inter-module SWAPs). By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
Parallel Translation and Optimization
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the module granularity (Figure 4(f)). Basis translation, in particular, is embarrassingly parallel: gates can be translated independently from each other. Meanwhile, as discussed in §2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural module boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
4. Methodology
We evaluate quantum compilers based on the quality of their solutions, their speed of compilation, and their ability to scale to large circuits/devices. To experimentally profile these characteristics, we benchmark a compiler on a suite of quantum circuits, detailed in §4.1, targeting a mock module-based quantum backend, as specified in §4.2. In total, we evaluate four compilers—a reference compiler, a peephole-augmented compiler, and two versions of SEQC—on simulated backends ranging from 1 to 100 modules. All compilers are implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.4 and Python $3.12.6$, and timing results are taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. For our reference compiler, we use the one built into Qiskit, parameterized with optimization_level=3. Because Qiskit is not natively module-aware, we run these experiments on a backend with universal inter-module links (§4.2). To construct the peephole compiler we augment the reference Qiskit compiler by inserting our peephole pass between its routing and basis_translation stages, thus making it module-aware. Our two SEQC versions differ in the technique they use for qubit-to-subcircuit mapping. Our first SEQC implementation leverages simulated annealing, initialized with a starting temperature $T_{0}=200$, a rate of change of $0.005T$ per round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core. Meanwhile, our second SEQC implementation leverages STA for subcircuit-to-qubit mapping and subcircuit-to-module mapping.
We aggregate results in a manner similar to SPEC benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits. The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of modules.
4.1. Benchmark Suite